# Algorithms and Lower Bounds for the Maximum Overlap of Two Polygons Under Translation

### v1

@@ 3 An ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{3/2}\log(nm))$ Algorithm
Lemma 3.1 ([10, 11]).
@@ 4.1 Construction
Anchor gadget
Both $P$ and $Q$ have an anchor gadget, whose sole purpose is to restrict the set of translations which may attain the maximum. The anchor gadget $P_{A}$ of $P$ is a square with side length $M$ including the anchor gadget $Q_{A}$ of $Q$ being a square of side length $1$ sufficiently far away from the origin. More precisely, the anchor gadget of $P$ is defined by the vertices $(0,100M)$, $(0,101M+1)$, $(M,101M+1)$ and $(M,100M)$, and the anchor gadget of $Q$ is defined by the vertices $(0,100M)$, $(0,100M+1)$, $(1,100M+1)$ and $(1,100M)$. Observe in particular that the anchor gadgets lie on the grid with side length $\color[rgb]{0,0,0}\mathrm{\varepsilon}$}}.
@@
Origin
Unlike the anchor gadget, which is far away from the origin, all other pieces of non-zero measure of $P$ and $Q$ are contained in $2M,3M]\times[-2M,3M]$, as we will see later. The rectangle $0,M]\times[0,M]$ is part of $P$ and serves as a root polygon to which its other gadgets are attached.
$x$-comb gadget
For the $\color[rgb]{0,0,0}i$}}th element $b_{{{\color[rgb]{0,0,0}i}}}\in B$ add the rectangle $P_{x}^{{{\color[rgb]{0,0,0}i}}}=[b_{{{\color[rgb]{0,0,0}i}}}+(3{{{\color[rgb]{0,0,0}i}}}+1){{\color[rgb]{0,0,0}\mathrm{\varepsilon}}},b_{{{\color[rgb]{0,0,0}i}}}+(3{{{\color[rgb]{0,0,0}i}}}+2){{\color[rgb]{0,0,0}\mathrm{\varepsilon}}}]\times[-2M,0]$ to $P$, which we refer to as the $\color[rgb]{0,0,0}i$}}}th $x$-prong of $P$. For the $\color[rgb]{0,0,0}k$}}}th element $d_{{{\color[rgb]{0,0,0}k}}}\in D$ we add the square $Q_{x}^{{{\color[rgb]{0,0,0}k}}}[-d_{{{\color[rgb]{0,0,0}k}}},-d_{{{\color[rgb]{0,0,0}k}}}+{{\color[rgb]{0,0,0}\mathrm{\varepsilon}}}]\times[-2M,-2M+{{\color[rgb]{0,0,0}\mathrm{\varepsilon}}}]$ to $Q$, which we refer to as the $\color[rgb]{0,0,0}k$}}}th $x$-prong of $Q$.
$y$-comb gadget
@@
sum gadget
For the $\color[rgb]{0,0,0}h$}}}th element $a_{{{\color[rgb]{0,0,0}h}}}\in A$ let $P_{d}^{{{\color[rgb]{0,0,0}h}}}$ be the polygon resulting from restricting the set $\{(x,y)\in{{\color[rgb]{0,0,0}\mathbb{R}}}^{2}\mid a_{{{\color[rgb]{0,0,0}h}}}\leq x+y\leq a_{{{\color[rgb]{0,0,0}h}}}+(10n+2){{\color[rgb]{0,0,0}\mathrm{\varepsilon}}}\}$ to $([-2M,M]\times[0,3M])\setminus([0,M]\times[0,M])$, and add it to $P$. We call the polygon corresponding to $a_{{{\color[rgb]{0,0,0}h}}}$ the $\color[rgb]{0,0,0}h$}}}th diagonal prong of $P$. Lastly add the single square $Q_{v}=[-2M,-2M+{{\color[rgb]{0,0,0}\mathrm{\varepsilon}}}]\times[2M,2M+{{\color[rgb]{0,0,0}\mathrm{\varepsilon}}}]$ to $Q$, called the verifier of $Q$.
Intuition
The role of the anchor gadgets is to restrict the set of translations $\tau$ such that ${{\color[rgb]{0,0,0}\mathrm{area}}}(P\cap(Q+\tau))>1$ to $0,M]\times[0,M]$. This is achieved since outside the anchor gadget of $Q$ (which has measure exactly $1$), $Q$ has a total measure of $(2n+1){{\color[rgb]{0,0,0}\mathrm{\varepsilon}}}\ll 1$. Hence, any translation that attains an overlap of more than $1$ must place the anchor gadget of $Q$ within the anchor gadget of $P$ and move at least one prong of $Q$ inside $P$.

### v2

@@ 3 An ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{3/2}\log(nm))$ Algorithm
Lemma 3.1 ([dobkin1985linear, dobkin1990separation]).
@@ 4.1 Construction
Anchor gadget: Both $P$ and $Q$ have an anchor gadget, whose sole purpose is to restrict the set of translations which may attain the maximum. The anchor gadget $P_{A}$ of $P$ is a square with side length $M$ including the anchor gadget $Q_{A}$ of $Q$ being a square of side length $1$ sufficiently far away from the origin. More precisely, the anchor gadget of $P$ is defined by the vertices $(0,100M)$, $(0,101M+1)$, $(M,101M+1)$ and $(M,100M)$, and the anchor gadget of $Q$ is defined by the vertices $(0,100M)$, $(0,100M+1)$, $(1,100M+1)$ and $(1,100M)$. Observe in particular that the anchor gadgets lie on the grid with side length $\color[rgb]{0,0,0}\mathrm{\varepsilon}$}}.
@@
Origin: Unlike the anchor gadget, which is far away from the origin, all other pieces of non-zero measure of $P$ and $Q$ are contained in $2M,3M]\times[-2M,3M]$, as we will see later. The rectangle $0,M]\times[0,M]$ is part of $P$ and serves as a root polygon to which its other gadgets are attached.
$x$-comb gadget: For the $\color[rgb]{0,0,0}i$}}th element $b_{{{\color[rgb]{0,0,0}i}}}\in B$ add the rectangle $P_{x}^{{{\color[rgb]{0,0,0}i}}}=[b_{{{\color[rgb]{0,0,0}i}}}+(3{{{\color[rgb]{0,0,0}i}}}+1){{\color[rgb]{0,0,0}\mathrm{\varepsilon}}},b_{{{\color[rgb]{0,0,0}i}}}+(3{{{\color[rgb]{0,0,0}i}}}+2){{\color[rgb]{0,0,0}\mathrm{\varepsilon}}}]\times[-2M,0]$ to $P$, which we refer to as the $\color[rgb]{0,0,0}i$}}}th $x$-prong of $P$. For the $\color[rgb]{0,0,0}k$}}}th element $d_{{{\color[rgb]{0,0,0}k}}}\in D$ we add the square $Q_{x}^{{{\color[rgb]{0,0,0}k}}}[-d_{{{\color[rgb]{0,0,0}k}}},-d_{{{\color[rgb]{0,0,0}k}}}+{{\color[rgb]{0,0,0}\mathrm{\varepsilon}}}]\times[-2M,-2M+{{\color[rgb]{0,0,0}\mathrm{\varepsilon}}}]$ to $Q$, which we refer to as the $\color[rgb]{0,0,0}k$}}}th $x$-prong of $Q$.
@@
sum gadget: For the $\color[rgb]{0,0,0}h$}}}th element $a_{{{\color[rgb]{0,0,0}h}}}\in A$ let $P_{d}^{{{\color[rgb]{0,0,0}h}}}$ be the polygon resulting from restricting the set $\{(x,y)\in{{\color[rgb]{0,0,0}\mathbb{R}}}^{2}\mid a_{{{\color[rgb]{0,0,0}h}}}\leq x+y\leq a_{{{\color[rgb]{0,0,0}h}}}+(10n+2){{\color[rgb]{0,0,0}\mathrm{\varepsilon}}}\}$ to $([-2M,M]\times[0,3M])\setminus([0,M]\times[0,M])$, and add it to $P$. We call the polygon corresponding to $a_{{{\color[rgb]{0,0,0}h}}}$ the $\color[rgb]{0,0,0}h$}}}th diagonal prong of $P$. Lastly add the single square $Q_{v}=[-2M,-2M+{{\color[rgb]{0,0,0}\mathrm{\varepsilon}}}]\times[2M,2M+{{\color[rgb]{0,0,0}\mathrm{\varepsilon}}}]$ to $Q$, called the verifier of $Q$.
Intuition: The role of the anchor gadgets is to restrict the set of translations $\tau$ such that ${{\color[rgb]{0,0,0}\mathrm{area}}}(P\cap(Q+\tau))>1$ to $0,M]\times[0,M]$. This is achieved since outside the anchor gadget of $Q$ (which has measure exactly $1$), $Q$ has a total measure of $(2n+1){{\color[rgb]{0,0,0}\mathrm{\varepsilon}}}\ll 1$. Hence, any translation that attains an overlap of more than $1$ must place the anchor gadget of $Q$ within the anchor gadget of $P$ and move at least one prong of $Q$ inside $P$.